# Investigating FRB 20240114A with FAST: Morphological Classification and Drifting Rate Measurements in a Burst-Cluster Framework

### v1

@@ Abstract
This study investigates the morphological classification and drifting rate measurement of the repeating fast radio burst (FRB) source FRB 20240114A using the Five-hundred-meter Aperture Spherical Telescope (FAST). Detected on January 14, 2024, FRB 20240114A exhibited an exceptionally high burst rate, revealing its unique properties. Through comprehensive observational campaigns spanning several months, we chose an extensive dataset, including 3,203 bursts (2,109 burst-clusters in total) during a continuous monitoring session (15,780 seconds) on March 12, 2024. Improving upon previous work, we clarify the definitions of sub-bursts, bursts and burst clusters. Using an average dispersion measure (DM) of 529.2 pc cm-3, we classified the burst-clusters into several categories, including Downward Drifting, Upward Drifting, No Drifting, No Evidence for Drifting, Not-Clear, and Complex burst-clusters. Notably, 233 burst-clusters (23.82%) exhibited upward drifting, which accounts for this proportion of the total 978 burst-clusters showing drifting behavior. Additionally, if 142 upward drifting single-component burst-clusters are excluded, upward drifting double- and multiple-components burst-clusters still account for 10.89% of the total sample of 836 burst-clusters exhibiting drifting behavior, equating to 91 burst-clusters. Furthermore, if only upward drifting burst-clusters with consecutive time intervals (or upward drifting bursts) are considered, only 9 bursts remain. Drifting rate comparisons with other physical quantities reveal a robust inverse correlation between the drifting rate and sub-burst effective widths. In single-component burst-clusters, those with upward drifting exhibit smaller effective widths, bandwidths, and fluxes compared to their downward drifting counterparts. Moreover, for upward drifting single-component burst-clusters, the drifting rate increases with peak frequency. A Kolmogorov-Smirnov test further indicates that upward drifting burst-clusters possess longer consecutive time intervals than downward drifting ones, suggesting distinct underlying physical mechanisms.
1 Introduction
@@
Fast radio bursts (FRBs) are brief radio flashes, typically lasting milliseconds (Lorimer et al., 2007). They exhibit high dispersion measures (DM), exceeding the maximum estimated from the electron column density model (Cordes & Lazio, 2003; Yao et al., 2017) for the Milky Way, making almost all of them considered extragalactic in origin (Thornton et al., 2013; Cordes & Chatterjee, 2019; Petroff et al., 2019; Zhang, 2020, 2023). Since the discovery of the first event, FRB 20010724 (Lorimer et al., 2007), hundreds of FRB sources have been discovered, with a small number of sources observed to emit repeated bursts. After the discovery of the first repeater FRB 20121102A (Spitler et al., 2014), a large number of repeated bursts have been detected from 69 FRB repeating sources [^1] https://blinkverse.zero2x.org/, up to March 2025 (CHIME/FRB Collaboration et al., 2019a; Fonseca et al., 2020; CHIME/FRB Collaboration et al., 2021; Chime/Frb Collaboration et al., 2023). These FRBs exhibit brightness temperatures ($T_{b}>10^{35}$ K), indicating that their emission mechanisms must be coherent (Zhang, 2023). With the advent of new-generation radio telescopes, such as the Canadian Hydrogen Intensity Mapping Experiment (CHIME [^2] https://chime-experiment.ca/en), the Green Bank Telescope (GBT [^3] https://greenbankobservatory.org/), the Australian Square Kilometer Array Pathfinder (ASKAP [^4] https://www.csiro.au/en/about/facilities-collections/ATNF/ASKAP-radio-telescope), the Karoo Array Telescope (MeerKAT [^5] https://www.sarao.ac.za/science/meerkat/), the Very Large Array (VLA [^6] https://public.nrao.edu/telescopes/VLA/), the Low-Frequency Array (LOFAR [^7] https://www.astron.nl/telescopes/lofar/), the Westerbork Synthesis Radio Telescope (WSRT [^8] https://www.astron.nl/telescopes/wsrt-apertif/), the Kunming 40-Meter Radio Telescope (KM40M [^9] https://ynao.cas.cn/kyzb/202307/t20230718_6812150.html) and the Shanghai Tianma Radio Telescope (TMRT [^10] http://65m.shao.cas.cn/) which offer higher time resolution and broader fields of view, the detection of FRBs has become more frequent and precise, further accelerating the development of this field. In particular, China’s Five-Hundred-Meter Aperture Spherical Radio Telescope (FAST [^11] https://fast.bao.ac.cn/) (Jiang et al., 2019, 2020; Qian et al., 2020), with its extremely high sensitivity and wide field of view, has become an important tool for FRB observations and has provided a wealth of valuable data.
@@
Although most FRBs still appear as one-off burst events, the study of repeating FRBs provides in-depth insight into repeated burst-clusters and drives the exploration of their potential origins (magnetars, black holes, or other extreme astrophysical environments). For further discussion on the historical development and physical mechanisms of FRB, refer to relevant reviews (Zhang, 2020, 2023; Cordes & Chatterjee, 2019; Petroff et al., 2019; Platts et al., 2019). In particular, observations of repeating FRBs have consistently reported a phenomenon often called the “sad trombone” effect. This pattern involves sub-bursts that appear at progressively lower frequencies over time. It was initially highlighted in FRB 20121102A (Hessels et al., 2019) and has since been documented in additional sources (CHIME/FRB Collaboration et al., 2019a; Chime/Frb Collaboration et al., 2020; Fonseca et al., 2020). Later studies confirmed that this downward drifting is widespread and can be detected at various observing frequencies (Hilmarsson et al., 2021; Jahns et al., 2023; Zhou et al., 2022; Zhang et al., 2022). Detailed analyses have shown that multiple sub-bursts can each shift to lower frequencies as they evolve (Platts et al., 2021; Pleunis et al., 2021), and recent work demonstrated that this effect extends to low radio bands of around a few hundred megahertz (Fonseca et al., 2020; Pastor-Marazuela et al., 2021). These phenomena have motivated diverse theoretical explanations.
@@
The apparent downward drifting feature can be readily interpreted within the framework magnetospheric models invoking the so-called radius-to-frequency mapping (Wang et al., 2019; Lyutikov, 2020; Zhang, 2022). In particular, if FRBs are produced by bunches via curvature radiation or inverse Compton scattering, the line of sight would usually sweep the emitters at lower altitudes earlier and those at higher altitudes later (Wang et al., 2019; Zhang, 2022; Qu & Zhang, 2024) so that lower-frequency emission arrives later. In order to interpret upward drifting features, one needs to invoke different injection times of bunches from different field lines Wang et al. (2020). Since the initial condition is more contrived, the chance of observing upward drifting is lower than that of observing downward drifting. Within far-away models that invoke synchrotron maser in relativistic shocks, the downward drifting feature can be accounted for if the shock undergoes deceleration so that there is continuous decay of the spectral break (Metzger et al., 2019, 2022). Independently, studies incorporating relativistic bulk motion and Dicke’s superradiance have demonstrated that such kinematic effects can predict inverse correlations between drifting rates and sub-burst durations (Rajabi et al., 2020).
@@
Although these emission mechanisms address the intrinsic properties of FRBs, propagation effects through the interstellar and circum-burst media introduce additional complexities: scintillation in turbulent media modulates burst morphology (Rickett, 1990), temporal variations in the DM of a repeating FRB could potentially indicate changes in the electron density along the propagation path or in the immediate vicinity of the source (Zhao & Wang, 2021), while plasma lensing from electron density gradients (Cordes et al., 2017) can mimic downward drifting via chromatic deflection. Tuntsov et al. (2021) proposed that a post-dispersion modulating screen imprints temporally sharp broadband structures, with subsequent dispersion in the circum-burst plasma generating the observed drift. Their model further posits that emission is beamed away from Earth, requiring chromatic deflection by plasma prisms ($\alpha\propto\nu^{-2}$) to redirect radiation into our line of sight, a mechanism that simultaneously explains the narrow observed bandwidths despite broadband intrinsic emission. However, the overwhelming prevalence of downward drifting events, combined with rare upward drifting bursts (Pleunis et al., 2021), suggests that a single physical framework may be insufficient to explain the full diversity of observed drifting phenomena.
@@
Temporal analyses across the FRB population reveal two key patterns: (1) bimodal waiting-time distributions in repeaters, suggesting distinct triggering regimes (Li et al., 2021) or potentially linked to starquake-aftershock sequences (Luo et al., 2025); (2) periodic activity windows like FRB 20180916B’s 16.35-day cycle, indicative of orbital/precessional modulation (Rajwade et al., 2020). Recently, a potential second-scale periodicity has been reported in the repeating FRB 20201124A, with periods on the order of 1.7 seconds detected on two separate observation days, which may suggest a possible local periodicity that could be associated with the rotation of the central engine, although no periodic signal was found in other datasets (Du et al., 2025). Notably absent are sub-second periodicities matching magnetar spin periods, with upper limits ($P<10$ ms) ruling out canonical pulsar-like beam sweeping (Niu et al., 2022; Du et al., 2024). For FRB 20240114A, comprehensive timing analysis based on FAST data is underway to search for potential periodicities (Dengke Zhou, 2025), a study motivated by its recent detection by CHIME, which revealed an exceptional burst rate ($>100\,\mathrm{hr}^{-1}$) and a high dispersion measure (DM = 527.7 pc cm-3) (Chime/Frb Collaboration et al., 2023; Shin & CHIME/FRB Collaboration, 2024; Shin et al., 2025), as well as by the observed near-complete linear polarization ($>95\%$) (Xie et al., 2024), a signature that is consistent with coherent curvature radiation in ordered magnetospheric fields (Wang et al., 2019). If the downward drifting and upward drifting are caused by DM misunderstanding, the corrected timing may strengthen the two pillars. Huang et al. (2025) report observations of the repeating FRB 20240114A using the KM40M and FAST, revealing a flattened high-energy tail in its burst energy distribution that aligns with properties of non-repeating FRBs, supporting a potential common origin between repeating and one-off FRBs, while Xie et al. (2025) demonstrate using the GBT that FRB 20240114A exhibits the highest burst rate ever recorded with (sub)-100-meter radio telescopes (264 hr-1) and extreme linear polarization ($>90\%$ in 81% of bursts).
@@
In this work, we focus solely on the data obtained from the FAST on 2024-03-12, which provided the longest-duration monitoring and the highest burst rate of FRB 20240114A observed to date. The observation lasted 15,780 seconds, during which a large number of bursts were detected. These abundant data allow us to investigate the properties of repeating bursts in considerable detail. This paper is organized as follows. Section 2 introduces the data preprocessing procedure and the calculation of various physical quantities. Section 3 focuses on the morphological classification, providing definitions for sub-burst, burst, and burst-cluster, along with the concepts of consecutive and intermittent time intervals, while Section 4 presents drifting rate measurements, discusses their correlation with other physical quantities, and highlights the distinct recurrence patterns between upward and downward drifting events, suggesting that multi-mechanism models may be necessary to fully explain FRB dynamics.
2 Observations and burst-cluster Detection
Following its discovery by CHIME on 2024 January 14, FRB 20240114A was subsequently targeted for intensive follow-up observations with the FAST. The observations were carried out using the center beam of the L-band 19-beam receiver (Jiang et al., 2020), initially pointed to R.A. = 21:27:39.888, decl. = +04:21:00.36, as reported by CHIME (Shin & CHIME/FRB Collaboration, 2024). On 2024 February 15, the pointing was calibrated to the position R.A. = 21:27:39.84, decl. = +04:19:46.3, localized by MeerKAT (Tian et al., 2024). After obtaining the new coordinates, the subsequent observations with FAST were also updated accordingly. The monitoring campaign spanned from 2024 January 28 to 2024 August 30, comprising 59 separate observing sessions and generating a cumulative on-source integration time of 34.7 hours. Data acquisition was performed in pulsar searching mode, with four polarization channels (XX, YY, XY*, and X*Y) sampled at a time resolution of 49.152 $\mu$s by the digital backend. A total of 4096 frequency channels, each with a bandwidth of 0.122070 MHz, were used to fully cover the 1.0–1.5 GHz range. All data were recorded in 8-bit PSRFITS format (Hotan et al., 2004), with each FITS file containing 6.442450944 s of data. The detection threshold $7\sigma$ of FAST was measured as $0.015~{}\mathrm{Jy~{}ms}$ for bursts with a width of $1~{}\mathrm{ms}$ (Li et al., 2021). This sensitivity represents at least a threefold improvement compared to previous Arecibo observations (Mathews, 2013), enabling the detection of fainter bursts.
@@ 2 Observations and burst-cluster Detection
These high time-resolution observations provide a rich data set for investigating the burst properties and temporal evolution of FRB 20240114A. At the beginning of each session, calibration signals consisting of a 10 K periodic noise were injected for one minute before each session to calibrate the system temperature ($T_{\rm sys}$). The temporal variation in background was taken into account by considering the telescope gain dependence on the zenith angle and frequency (Jiang et al., 2020), enabling us to convert the average peak flux from Kelvin to Jansky units. We then computed each burst isotropic-equivalent energy via
@@
where $D_{L}\approx 630.72$ Mpc and $z=0.13$ for FRB 20240114A (Bhardwaj et al., 2024).
For additional details on the calibration procedures and a comprehensive overview of all 59 observing sessions, including the data collected during the March 12 2024 session, we refer the reader to our first paper in this series (Jun-Shuo Zhang, 2025). In the present work, we focus solely on the data collected during the March 12 2024 session, with the burst timestamps obtained from the analysis pipeline described in that paper.
2.1 Data Analysis and Burst Detection

### v3

@@ Abstract
This study investigates the morphological classification and drifting rate measurement of the repeating fast radio burst (FRB) source FRB 20240114A using the Five-hundred-meter Aperture Spherical Telescope (FAST). Detected on January 14, 2024, FRB 20240114A exhibited an exceptionally high burst rate, revealing unique properties. Through observational campaigns over several months, we selected a dataset comprising 3,203 bursts (2,109 burst-clusters) during a continuous monitoring session (15,780 seconds) on March 12, 2024. Improving upon previous work, we clarify the definitions of sub-bursts, bursts and burst-clusters. Using an average dispersion measures (DM) of 529.2 pc cm-3, we classified the burst-clusters into Downward Drifting, Upward Drifting, No Drifting, No Evidence for Drifting, Not-Clear, and Complex burst-clusters. Among the 978 burst-clusters that exhibit drifting behavior, 233 (23.82%) show upward drifting. Additionally, if 142 upward drifting single-component burst-clusters are excluded, upward drifting double- and multi-component burst-clusters still account for 10.89% of the 836 burst-clusters exhibiting drifting behavior, equating to 91 burst-clusters. Furthermore, if only upward drifting burst-clusters with consecutive time intervals (or upward drifting bursts) are considered, only 9 bursts remain. Drifting rate comparisons with other physical quantities reveal that the drifting rate increases with peak frequency for single-component burst-clusters with drifting behavior. Moreover, in single-component burst-clusters, those with upward drifting exhibit smaller effective widths, bandwidths, and fluxes than their downward drifting counterparts. A Kolmogorov-Smirnov test further indicates that upward drifting burst-clusters possess longer consecutive time intervals than downward drifting ones, suggesting distinct underlying physical mechanisms.
I Introduction
@@
Fast radio bursts (FRBs) are brief radio flashes, typically lasting milliseconds (2007Sci...318.. [^7] L). They exhibit high dispersion measures (DMs), exceeding the maximum estimated from the electron column density model (cordes2003ne2001inewmodelgalactic; 2017ApJ...835...29Y) for the Milky Way, making almost all of them considered extragalactic in origin (doi:10.1126/science.1236789; annurev-astro-091918-104501; 2019A&ARv..27....4P; 2020Natur.587...45Z; 2023RvMP...95c5005Z). Since the discovery of the first event, FRB 20010724 (2007Sci...318.. [^7] L), hundreds of FRB sources have been discovered, with a small number of sources observed emitting repeated bursts. After the discovery of the first repeater, FRB 20121102A (2014ApJ...790..101S), a large number of repeated bursts have been detected from 72 repeating FRB sources [^1] https://blinkverse.zero2x.org/, up to September 2025 (2019ApJ...885L..24C; 2020ApJ...891L...6F; 2021ApJS..257...59C; 2023ApJ...947...83C). These FRBs exhibit high brightness temperatures ($T_{b}>10^{35}$ K), indicating that their emission mechanisms must be coherent (2023RvMP...95c5005Z). With the advent of new-generation radio telescopes, such as the Canadian Hydrogen Intensity Mapping Experiment (CHIME; 2018ApJ...863...48C 2018ApJ...863...48C), the Green Bank Telescope (GBT; 2020AAS...23624305W 2020AAS...23624305W), the Australian Square Kilometer Array Pathfinder (ASKAP; 2022arXiv220808245K 2022arXiv220808245K), the Karoo Array Telescope (MeerKAT; 2016mks..confE...1J 2016mks..confE...1J), the Very Large Array (VLA; 2011ApJ...739L...1P 2011ApJ...739L...1P), the Low-Frequency Array (LOFAR; 2013A&A...556A...2V 2013A&A...556A...2V), the Westerbork Synthesis Radio Telescope (WSRT; 2022A&A...658A.146V 2022A&A...658A.146V), the Kunming 40-Meter Radio Telescope (KM40M; 2010RAA....10..805H 2010RAA....10..805H) and the Shanghai Tianma Radio Telescope (TMRT; 2024AstTI...1..239L 2024AstTI...1..239L) which offer higher time resolution and broader fields of view, the detection of FRBs has become more frequent and precise, further accelerating the development of this field. In particular, China’s Five-Hundred-Meter Aperture Spherical Radio Telescope (FAST; 2011IJMPD..20..989N; 2019SCPMA..6259502J; 2020Innov...100053Q), with its extremely high sensitivity, has become an important tool for FRB observations and has provided a wealth of valuable data.
@@
Although most FRBs still appear as one-off burst events, the study of repeating FRBs provides in-depth insight into repeated burst-clusters and drives the exploration of their potential origins (magnetars, black holes, or other extreme astrophysical environments). For further discussion on the historical development and physical mechanisms of FRBs, refer to relevant reviews (2020Natur.587...45Z; 2023RvMP...95c5005Z; 2019ARA&A..57..417C; 2019A&ARv..27....4P; 2019PhR...821....1P). In particular, observations of repeating FRBs have consistently reported a phenomenon often called the “sad trombone” effect. This pattern involves sub-bursts that appear at progressively lower frequencies over time. It was initially highlighted in FRB 20121102A (Hessels_2019) and has since been documented in additional sources (2019ApJ...885L..24C; 2020Natur.582..351C; 2020ApJ...891L...6F). Later studies confirmed that this downward drifting is widespread and can be detected at various observing frequencies (2021MNRAS.508.5354H; 2023MNRAS.519.. [^6] J; 2022RAA....22l4001Z; 2022RAA....22l4002Z). Detailed analyses have shown that multiple sub-bursts can each shift to lower frequencies as they evolve (2021MNRAS.505.3041P; 2021ApJ...923....1P), and recent work demonstrated that this effect extends to low radio bands of around a few hundred megahertz (2020ApJ...891L...6F; 2021Natur.596..505P). These phenomena have motivated diverse theoretical explanations, which are elaborated in the following discussion.
@@
The apparent downward drifting feature can be readily interpreted within the framework of magnetospheric models invoking the so-called radius-to-frequency mapping (2019ApJ...876L..15W; osti_1802960; 2022ApJ...925...53Z). In particular, if FRBs are produced by bunches via curvature radiation or inverse Compton scattering, the line of sight would usually sweep the emitters at lower altitudes earlier and those at higher altitudes later (2019ApJ...876L..15W; 2022ApJ...925...53Z; 2024ApJ...972..124Q) so that lower-frequency emission arrives later. In order to interpret upward drifting features, one needs to invoke different injection times of bunches from different field lines (2020ApJ...899..109W). This is a less natural assumption, and therefore, such events are expected to be less common than downward drifting. Within far-away models that invoke the synchrotron maser in relativistic shocks, the downward drifting feature can be accounted for if the shock undergoes deceleration so that there is continuous decay of the spectral break (10.1093/mnras/stz700; 2022ApJ...925..135M). Independently, studies incorporating relativistic bulk motion and Dicke’s superradiance have demonstrated that such kinematic effects can predict inverse correlations between drifting rates and sub-burst durations (2020MNRAS.498.4936R).
@@
Although these emission mechanisms address the intrinsic properties of FRB, propagation effects through the interstellar and circum-burst media introduce additional complexities: scintillation in turbulent media modulates burst morphology (1990ARA&A..28..561R); temporal variations in the DM of a repeating FRB could potentially indicate changes in the electron density along the propagation path or in the immediate vicinity of the source (Zhao:2021vns); while plasma lensing from electron density gradients (2017ApJ...842...35C) can mimic downward drifting via chromatic deflection. 2021arXiv210713549T proposed that a post-dispersion modulating screen imprints temporally sharp broadband structures, with subsequent dispersion in the circum-burst plasma generating the observed drift. Their model further posits that emission is beamed away from Earth, requiring chromatic deflection by plasma prisms ($\alpha\propto\nu^{-2}$) to redirect radiation into our line of sight, a mechanism that simultaneously explains the narrow observed bandwidths despite the broadband intrinsic emission. However, the overwhelming prevalence of downward drifting events, combined with rare upward drifting bursts (2021ApJ...923....1P), suggests that a single physical framework may be insufficient to explain the full diversity of the observed drifting phenomena.
@@
Temporal analyses across the FRB population reveal two key patterns: (1) bimodal waiting-time distributions in repeaters, suggesting distinct triggering regimes (2021Natur.598..267L) or potentially linked to starquake-aftershock sequences (2025arXiv250216626L; 2025arXiv250812567Q); (2) periodic activity windows, such as FRB 20180916B’s 16.35-day cycle, indicative of orbital/precessional modulation (2020Natur.582..351C). Recently, a potential second-scale periodicity has been reported in the repeating FRB 20201124A, with periods on the order of 1.7 seconds detected on two separate observation days. This may suggest a possible intrinsic periodicity associated with the rotation of the central engine, in contrast to longer-term orbital or precessional cycles; although no periodic signal was found in other datasets (du2025secondscaleperiodicityactiverepeating). Notably absent are sub-second periodicities matching magnetar spin periods, with upper limits ($P<10$ ms) ruling out canonical pulsar-like beam sweeping (2022RAA....22l4004N; 2024ApJ...977..129D). For FRB 20240114A, comprehensive timing analysis based on FAST data is underway to search for potential periodicities (2025arXiv250714708Z), a study motivated by its recent detection by CHIME, which revealed an exceptional burst rate ($100\,\mathrm{hr}^{-1}$) and a high dispersion measure (DM = 527.7 pc cm-3) (2023ApJ...947...83C; 2024ATel16420....1S; 2025arXiv250513297S), as well as the observed near-complete linear polarization ($95\%$) (2024arXiv241010172X), a signature that is consistent with coherent curvature radiation in ordered magnetospheric fields (2019ApJ...876L..15W). If the downward drifting and upward drifting are partly caused by incorrect DMs, proper dedispersion correction would clarify these patterns and strengthen the two temporal pillars outlined above. 2025arXiv250403569H reported observations of the repeating FRB 20240114A using the KM40M and FAST, revealing a flattened high-energy tail in its burst energy distribution that aligns with the properties of non-repeating FRBs, supporting a potential common origin between repeating and one-off FRBs, while 2025ApJS..278...49X demonstrated using the GBT that FRB 20240114A exhibits the highest burst rate ever recorded with (sub)-100-meter radio telescopes (264 hr-1) and extreme linear polarization ($90\%$ in 81% of bursts). In addition, long-term simultaneous monitoring at 2.25/8.60 GHz with TMRT detected 155 bursts at 2.25 GHz but none at 8.60 GHz, placing strong constraints on the high-frequency activity of FRB 20240114A (2025ApJ...992..185W).
@@
In this work, we focus solely on the data obtained from the FAST on March 12, 2024, which provided the longest-duration monitoring and the highest burst rate of FRB 20240114A observed to date. The observation lasted 15,780 seconds, during which a large number of bursts were detected. These abundant data allow us to investigate the properties of repeating bursts in considerable detail. This paper is organized as follows. Section II introduces the data preprocessing procedure and the calculation of various physical quantities. Section III focuses on the morphological classification, providing definitions for sub-burst, burst, and burst-cluster, along with the concepts of consecutive and intermittent time intervals. Section IV presents drifting rate measurements, discusses their correlation with other physical quantities, and highlights the distinct recurrence patterns between upward and downward drifting events, suggesting that multi-mechanism models may be necessary to fully understand FRB mechanisms.
II Observations and burst-cluster Detection
[…]
@@ II Observations and burst-cluster Detection
These high time-resolution observations provide a rich dataset for investigating the burst properties and temporal evolution of FRB 20240114A. At the beginning of each session, calibration signals consisting of a 10 K periodic noise were injected for one minute before each session to calibrate the system temperature ($T_{\rm sys}$). The temporal variation in background was taken into account by considering the telescope gain dependence on the zenith angle and frequency (2020RAA....20...64J), enabling us to convert the antenna temperature (K) to flux density (Jy). We then computed each burst isotropic-equivalent energy via
@@
where $D_{L}\approx 630.72$ Mpc and $z=0.13$ for FRB 20240114A (2024ATel16613....1B).
For additional details on the calibration procedures and a comprehensive overview of all 59 observing sessions, including the data collected during the March 12, 2024 session, we refer the reader to our first paper in this series (2025arXiv250714707Z). In the present work, we focus solely on the data collected during the March 12, 2024 session, with the burst timestamps obtained from the analysis pipeline described in that paper.
II.1 Data Analysis and Burst Detection